### v2

@@ Abstract
Recent transportation research highlights the potential of autonomous vehicles (AV) to improve traffic flow mobility as they are able to maintain smaller car-following distances. However, as a unique class of ground robots, AVs are susceptible to robotic errors, particularly in their perception and control modules, leading to uncertainties in their movements and an increased risk of collisions. Consequently, conservative operational strategies, such as larger headway and slower speeds, are implemented to prioritize safety over mobility in real-world operations. To reconcile the inconsistency, this paper presents an analytical model framework that delineates the endogenous reciprocity between traffic safety and mobility that arises from AVs’ robotic uncertainties. Using both realistic car-following data and a stochastic intelligent driving model (IDM), the stochastic car-following distance is derived as a key parameter, enabling analysis of single-lane capacity and the collision probability. A semi-Markov process is then employed to model the dynamics of the lane capacity, and the resulting collision-inclusive capacity, representing expected lane capacity under stationary conditions, serves as the primary performance metric for fully autonomous traffic. The analytical results are further utilized to investigate the impacts of critical parameters in AV and roadway designs on traffic performance, as well as the properties of optimal speed and headway under mobility-targeted or safety-dominated management objectives. Extensions to scenarios involving multiple non-independent collisions or multi-lane traffic scenarios are also discussed, which demonstrates the robustness of the theoretical results and their practical applications.
@@ 1 Introduction
The emergence of autonomous vehicles (AVs), or fully automated vehicles, offers the potential to improve the overall safety and mobility of road transportation (Fernandes and Nunes, 2012; Jiménez et al., 2016). AVs with intelligent decision-making abilities and accurate machinery operations will be able to prevent collisions caused by human misbehavior, which is identified as the leading cause of car crashes (Mueller et al., 2020). When investigating their impact on traffic mobility, AVs are usually treated as ideal machines with more advanced driving capabilities. Accordingly, they are assumed to maintain a shorter stable headway in traffic streams than human-driven vehicles (HDVs) (Morando et al., 2018), contributing to different levels of increased roadway capacity (Ran and Tsao, 1996) as per their market penetration (Talebpour and Mahmassani, 2016; Chen et al., 2017; Seo and Asakura, 2017)  [^1] Seo and Asakura (2017) has shown AVs could improve the roadway capacity when their proportion in the mixed HDV-AV traffic flow reaches beyond some threshold..
@@
However, AVs’ robotic nature may prevent these benefits from being realized (Li et al., 2022). Like other ground robots, AV operation is streamlined into a standard process that consists of four modules: perception, localization, planning (or decision-making), and control (Levinson et al., 2011). While the first two modules rely on sensors that determine how accurately the AVs perceive their relative positional relationship with their surroundings, the latter two represent their level of driving intelligence. Each module contains a certain level of uncertainty, which together cause a chance of system error, leading an AV a deviation from its ideal movement or even a collision with the surrounding objects. Adopting conservative driving strategies, such as slower speed and longer car-following distance, could alleviate the negative impact of AVs’ systematic uncertainty on driving safety performance (Sybis et al., 2020). As a trade-off, it would limit the traffic capacity and undermine the overall traffic mobility.
@@
Research efforts have been made to overcome the vehicular robotic deficiency, by reducing perception errors, improving localization accuracy (Wen et al., 2022), and developing robust control technologies against uncertainty (Huang et al., 2024), especially in Adaptive Cruise Control (ACC) or Cooperative Adaptive Cruise Control (CACC) (Zhou et al., 2017; Zhou and Ahn, 2019; Kontar and Ahn, 2022). Nevertheless, the system uncertainty could only be mitigated instead of being completely diminished (Mohammadian et al., 2023). In this respect, a growing research direction focuses on the verification and validation of AV safety performance under both simulated and realistic naturalistic driving environments (Waymo, 2017; Motors, 2018; Feng et al., 2021; Yan et al., 2023; Xu et al., 2022; Feng et al., 2023; Araujo et al., 2023), aiming to identify the critical scenarios that lead to AV safety hazard. Also closely intertwined with uncertainty, stability emerges as another key concern. From the empirical experiments of ACC systems in commercial vehicles, Makridis et al. (2021) and Li et al. (2021) identified the characteristics of oscillation amplification. Other studies, such as Zhang et al. (2020) and Wang et al. (2020), theoretically delineated the impact of uncertainty, particularly the positioning errors caused by robotic characteristics, on stability in automated car following systems. Traffic mobility, unfortunately, is usually overlooked in these studies. In fact, due to the absence of laws and regulations on AV traffic mobility (Shladover and Nowakowski, 2019; Vignon and Bahrami, 2025) and public concerns about autonomous driving collisions (Kyriakidis et al., 2015; Howard and Dai, 2014; Xing et al., 2022), autonomous driving companies usually adopt conservative driving approaches in real-road AV driving tests to achieve error-free safety performance. And as indicated earlier, when driving slower and keeping longer car-following distances than the surroundings, those pilot AVs naturally become moving bottlenecks that hold up regular traffic flows, generating traffic congestion and potential danger to roadways (Knoop et al., 2019; Schakel et al., 2017; McCarthy, 2022).
@@
This paper therefore aims to investigate the mutual relationship between traffic mobility and safety performance in a fully autonomous vehicle environment (e.g., a dedicated lane on a highway). Examining this hypothetical setting offers a long-term vision of potential challenges and benefits in autonomous driving systems, enabling proactive decision-making in AV development and regulation constructions at the current stage. Previous literature has only explored the trade-off between AV safety and mobility under uncertainty in two recent studies (Shi and Li, 2021; Li, 2022). While these studies reveal the stochastic behavior of autonomous vehicles and analyze their impact on macroscopic traffic from the angle of microscopic AV control, this study further contributes from two perspectives. First, an analytical model is developed to explicitly incorporate the uncertainties arising from AV operation process, providing a structured approach to mathematically relate safety and mobility performances. Second, macroscopic traffic flow considerations are emphasized beyond microscopic vehicle controls, seeking to provide insightful suggestions to AV manufacturers and traffic management authorities to ensure AV development aligns with societal benefits.
@@ 2 Robotic Uncertainty and Collision-Inclusive Capacity
Perception error is regarded as the vital error source that significantly impacts the motion of autonomous vehicles (Liu and Park, 2021). Perception information is usually modeled as random variables with specific distributions, inducing to a probability of deviation between observation and reality. When the deviated observation is fed into the subsequent planning and control modules, the error propagates through the AV operation process, resulting in a stochastic deviation between the actual motion and the expectation. In addition, the controller is physically limited and cannot achieve the planned motion perfectly and without delay. Moreover, due to the complex vehicle dynamics and the large amount of randomness present in real-world operations, such as slope, wind, road surface defects, etc., actuators can also generate certain errors when being controlled by autonomous driving (Zhou et al., 2017; Hu et al., 2022).
We refer to these perception errors, control errors, etc. as robotic uncertainties, as they are a problem that all autonomous robot systems will encounter. Therefore, even employing a theoretically collision-free driving strategy, AVs with robotic uncertainties may still have a chance to encounter collisions or any other unexpected behaviors. This characteristic and its impact are often ignored in the macroscopic traffic analysis of autonomous vehicles.
@@
The ego vehicle’s planning and control could be integrated as a function $\Gamma(\cdot)$, which maps the observed relative positions of surrounding object $i,\>i=1,2,...n$ being expressed by $X^{i}_{k}$ to its stochastic action $A_{e}^{k}$. Note that the function itself is deterministic once the input is given, but the actual motion is affected by control errors $O(A^{e}_{k})$:
$A^{e}_{k}=\Gamma(X^{1}_{k},X^{2}_{k},...,X^{n}_{k})+O(A^{e}_{k}).$ (1b)
@@ 3.2 Simulation of a high-order car-following model
Figure 2: Simulated stochastic car-following motion under robotic uncertainties based on IDM.
@@
Figure 3: A typical simulation result: (a) Acceleration of the ego vehicle and distance between two vehicles over time (b) Histogram of distance between two vehicles.
@@
$\epsilon_{d}\sim$ $\mathcal{N}(0,0.5)$ $\mathcal{N}(0,1)$ $\mathcal{N}(0,2)$
$\epsilon_{\Delta v}\sim$ $\mathcal{N}(0,0.5)$ $\mathcal{N}(0,1)$ $\mathcal{N}(0,2)$ $\mathcal{N}(0,0.5)$ $\mathcal{N}(0,1)$ $\mathcal{N}(0,2)$ $\mathcal{N}(0,0.5)$ $\mathcal{N}(0,1)$ $\mathcal{N}(0,2)$
$\epsilon_{\dot{v}}\sim\mathcal{N}(0,0.5)$ 0.0280 0.0236 0.0216 0.0280 0.0237 0.0222 0.0269 0.0234 0.0215
$\epsilon_{\dot{v}}\sim\mathcal{N}(0,1)$ 0.0478 0.0392 0.0324 0.0465 0.0388 0.0324 0.0449 0.0388 0.0317
$\epsilon_{\dot{v}}\sim\mathcal{N}(0,2)$ 0.0765 0.0688 0.0596 0.0756 0.0677 0.0582 0.0743 0.0665 0.0566
To test the robustness of the fitted Gaussian under different uncertainty conditions, we assume that the observation and control errors $\epsilon_{d}$, $\epsilon_{\Delta v}$, and $\epsilon_{\dot{v}}$ have variances of $0.5$, $1$, and $2$, while their mean values remain at 00. This results in a total of 27 fitness results, summarized in Table 1. The results indicate that NRMSEs are less than 0.08 regardless of the uncertainty conditions. Thus, it is reasonable to assume that the stochastic car-following distance follows a Gaussian distribution in a macroscopic statistical sense, and this assumption will be used for subsequent analysis.
@@ 3.3 Empirical AV data analysis
Considering that simulation experiments may not fully capture realistic driving behaviors, we also utilized real-world car-following data from commercial automated vehicles to explore their stochastic characteristics in longitudinal movements. The Waymo Open Dataset (Ettinger et al., 2021) was selected as the primary dataset for the empirical analysis. Zhou et al. (2024) processed the dataset, focusing on longitudinal car-following scenarios to accurately reflect the intricate behaviors of AVs. Key data representing the states of the leading vehicle (LV) and the following AV (FAV), including spatial gaps, speeds, accelerations, and positions, are recorded. Our previous study (Zhou et al., 2024) identified that Waymo’s data collection methods make it the only dataset capable of effectively capturing speed fluctuations in AV operations while maintaining stability. Therefore, it is well-suited for this study in identifying reliable results of stochastic car-following distances.
To identify stable car-following scenarios from complex real-world AV driving data, we conducted a sequence of data filtering steps. First, we selected scenarios where the preceding vehicle in front of the WayMo car maintained a steady speed, while the Waymo car tries to keep a consistent distance. Specifically, we excluded cases where the speed difference between two vehicles exceeded $1$ m/s, as larger differences could lead to a notable increase in car-following distance over a short period of time. Next, we filtered the cases to retain those where the leading vehicle maintaining an average speed of $20.2m/s$, the most common speed observed for the leading vehicles. This step controls the potential impact of speed differences on car-following distances, allowing us to focus purely on the intrinsic variations in AVs’ car-following behaviors. Finally, since the default car-following distance setting could vary across driving scenarios, we normalized the the car-following distance in all sampled data to have a uniform variance of one and a mean of zero, ensuring comparability across the dataset.
Figure 4: A real car-following data result: (a) Acceleration of FAV and distance between two vehicles over time (b) Histogram of distance between two vehicles.

### v3

@@ Abstract
Recent transportation research highlights the potential of autonomous vehicles (AV) to improve traffic flow mobility as they are able to maintain smaller car-following distances. However, as a unique class of ground robots, AVs are susceptible to robotic errors, particularly in their perception and control modules with imperfect sensors and actuators, leading to uncertainties in their movements and an increased risk of collisions. Consequently, conservative operational strategies, such as larger headway and slower speeds, are implemented to prioritize safety over mobility in real-world operations. To reconcile the inconsistency, this paper presents an analytical model framework that delineates the endogenous reciprocity between traffic safety and mobility that arises from AVs’ robotic uncertainties. Using both realistic car-following data and a stochastic intelligent driving model (IDM), the stochastic car-following distance is derived as a key parameter, enabling analysis of single-lane capacity and collision probability. A semi-Markov process is then employed to model the dynamics of the lane capacity, and the resulting collision-inclusive capacity, representing expected lane capacity under stationary conditions, serves as the primary performance metric for fully autonomous traffic. The analytical results are further utilized to investigate the impacts of critical parameters in AV and roadway designs on traffic performance, as well as the properties of optimal speed and headway under mobility-targeted or safety-dominated management objectives. Extensions to scenarios involving multiple non-independent collisions or multi-lane traffic scenarios are also discussed, which demonstrates the robustness of the theoretical results and their practical applications.
@@ 1 Introduction
The emergence of autonomous vehicles (AVs), or fully automated vehicles, offers the potential to improve the overall safety and mobility of road transportation (Fernandes and Nunes, 2012; Jiménez et al., 2016). AVs with intelligent decision-making abilities and accurate machinery operations will be able to prevent collisions caused by human misbehavior, which is identified as the leading cause of car crashes (Mueller et al., 2020). When investigating their impact on traffic mobility, AVs are usually treated as ideal machines with more advanced driving capabilities. Accordingly, they are assumed to maintain a shorter stable headway in traffic streams than human-driven vehicles (HDVs) (Morando et al., 2018), contributing to different levels of increased roadway capacity (Ran and Tsao, 1996) as per their market penetration (Talebpour and Mahmassani, 2016; Chen et al., 2017; Seo and Asakura, 2017)  [^1] Seo and Asakura (2017) has shown AVs could improve the roadway capacity when their proportion reaches beyond some threshold..
@@
However, AVs’ robotic nature may prevent these benefits from being realized (Li et al., 2022). Like other mobile robots, AV operation is streamlined into a standard process that consists of four modules: perception, localization, planning (or decision-making), and control (Levinson et al., 2011). While the first two modules rely on sensors that determine how accurately the AVs perceive their relative positional relationship with their surroundings, the latter two represent their level of driving intelligence. Each module contains a certain level of uncertainty, which together cause a chance of system error, leading an AV a deviation from its ideal movement or even a collision with the surrounding objects. We refer to these inherent errors as robotic uncertainties, as they are problems that all autonomous robot systems will encounter, but have nothing to do with unknown external environments or intentions. Adopting conservative driving strategies, such as slower speed and longer car-following distance, could alleviate the negative impact of AVs’ systematic uncertainty on driving safety performance (Sybis et al., 2020). As a trade-off, it would limit the traffic capacity and undermine the overall traffic mobility.
@@
Research efforts have been made to overcome the vehicular robotic deficiency, by reducing perception errors, improving localization accuracy (Wen et al., 2022), and developing robust control technologies against uncertainty (Huang et al., 2024), especially in Adaptive Cruise Control (ACC) or Cooperative Adaptive Cruise Control (CACC) (Zhou et al., 2017; Zhou and Ahn, 2019; Kontar and Ahn, 2022). Nevertheless, the system uncertainty could only be mitigated instead of being completely diminished (Mohammadian et al., 2023). In this respect, a growing research direction focuses on the verification and validation of AV safety performance under both simulated and realistic naturalistic driving environments (Waymo, 2017; Motors, 2018; Feng et al., 2021; Yan et al., 2023; Xu et al., 2022; Feng et al., 2023; Araujo et al., 2023), aiming to identify the critical scenarios that lead to AV safety hazard. Also closely intertwined with uncertainty, stability emerges as another key concern. From the empirical experiments of ACC systems in commercial vehicles, Makridis et al. (2021) and Li et al. (2021) identified the characteristics of oscillation amplification. Qin et al. (2023) further combine AV with HDV and study the stability and management of platoons under mixed traffic conditions. Other studies, such as Zhang et al. (2020) and Wang et al. (2020), theoretically delineated the impact of uncertainty, particularly the positioning errors caused by robotic characteristics, on stability in automated car following systems. Traffic mobility, unfortunately, is usually overlooked in these studies. In fact, due to the absence of laws and regulations on AV traffic mobility (Shladover and Nowakowski, 2019; Vignon and Bahrami, 2025) and public concerns about autonomous driving collisions (Kyriakidis et al., 2015; Howard and Dai, 2014; Xing et al., 2022), autonomous driving companies usually adopt conservative driving approaches in real-road AV driving tests to achieve error-free safety performance. And as indicated earlier, when driving slower and keeping longer car-following distances than the surroundings, those pilot AVs naturally become moving bottlenecks that hold up regular traffic flows, generating traffic congestion and potential danger to roadways (Knoop et al., 2019; Schakel et al., 2017; McCarthy, 2022).
@@
This paper therefore aims to investigate the mutual relationship between traffic mobility and safety performance in a fully autonomous vehicle environment, e.g., a dedicated lane on a highway. In the literature, two recent studies have also highlighted the trade-off between AV safety and mobility under uncertainty (Shi and Li, 2021; Li, 2022). Shi and Li (2021) validate the presence of stochastic behavior of autonomous vehicles from empirical vehicle trajectory data under different headway settings. Li (2022) further delineate the trade-off through an analytical approach grounded in a mathematical model of vehicle control dynamics, which also captures the relationship between stability and the two focal aspects: safety and mobility. This study further contributes from two perspectives. First, an analytical model is developed to explicitly incorporate the common robotic uncertainties arising from AV operation process, providing a structured approach to mathematically relate safety and mobility performances, which can be applied to both microscopic control models and empirical data. Second, macroscopic traffic flow considerations are emphasized beyond microscopic vehicle controls, connecting individual AV behavior to roadway-level outcomes, seeking to provide insightful suggestions to AV manufacturers and traffic management authorities to ensure AV development aligns with societal benefits.
@@ 2 Robotic Uncertainty and Collision-Inclusive Capacity
Perception error is regarded as the vital error source that significantly impacts the motion of autonomous vehicles (Liu and Park, 2021). Perception information is usually modeled as random variables with specific distributions, inducing to a probability of deviation between observation and reality. When the deviated observation is fed into the subsequent modules, the error propagates through the AV operation process, resulting in a stochastic deviation between the actual motion and the expectation. In addition, the controller is physically limited and cannot achieve the planned motion perfectly and without delay. Moreover, due to the complex vehicle dynamics and the large amount of randomness present in real-world operations, actuators can also generate certain errors when being controlled by autonomous driving (Zhou et al., 2017; Hu et al., 2022). These robotic uncertainties are inherent in all mobile robots with sensors, onboard algorithms, and actuators, even in an environment within operational design domain following established intentions. Therefore, even employing a theoretically collision-free driving strategy, AVs with robotic uncertainties may still have a chance to encounter collisions or any other unexpected behaviors. This characteristic and its impact are often ignored in the macroscopic traffic analysis of autonomous vehicles.
@@
The ego vehicle’s planning could be integrated as a function $\Gamma(\cdot)$, which maps the observed historical and current relative positions of surrounding object $i,\>i=1,2,...n$ being expressed by $X^{1:n}_{k-m_{h}:k}$ to its planning trajectory $F_{k:k+m_{f}}^{e}$:
$F_{k:k+m_{f}}^{e}=\Gamma(X^{1:n}_{k-m_{h}:k}).$ (1b)
The observed positions from $m_{h}$ steps ago ($k-m_{h}$) to the current step $k$ are utilized to optimize the perception results, make predictions of surrounding objects, and ultimately influence decision-making result. The planning trajectory $F_{k:k+m_{f}}^{e}$ looks $m_{f}$ steps ahead so as to generate a motion trajectory from current step $k$ to future ($k+m_{f}$). Note that the function itself is deterministic once the input is given, but the actual action $A_{e}^{k}$ is formed by the motion planning at the current time step $F_{k}^{e}$, combined with the control error $O(F_{k}^{e})$:
$A^{e}_{k}=F_{k}^{e}+O(F_{k}^{e}).$ (1c)
@@
Note that $H$ denotes the duration over which the transportation system is being analyzed. Since we focus on homogeneous autonomous vehicle streams with stability strategically ensured, the per-time-step collision probability of a single vehicle, $p$, is assumed to be independent of the choice of $H$.
@@ 3.2 Simulation of a high-order car-following model
Figure 2: Simulated stochastic car-following motion under robotic uncertainties based on IDM car-following model. The dynamics are simulated by MATLAB Simulink, and robotic errors are introduced into the perception of distance, the perception of speed, control.
@@
Figure 3: A typical simulation result: (a) Acceleration of the ego vehicle (red dotted line) and distance between two vehicles (blue circled line) over time; (b) Histogram of distance between two vehicles (blue bars) and the fitted Gaussian distribution (red line).
@@
To test the robustness of the fitted Gaussian under different uncertainty conditions, we assume that the observation and control errors $\epsilon_{d}$, $\epsilon_{\Delta v}$, and $\epsilon_{\dot{v}}$ have variances of $0.1$, $0.5$, and $1$, while their mean values remain at 00. The typical values represent the universal applicability of the perception capability of AVs as they cover precision from upper limit (such as the state-of-the-art performance in the NuScenes perception benchmark; Caesar et al., 2019) to lower limit (sub-meter level precision of millimeter wave radars). This results in a total of 27 fitness results, summarized in Table 1. The results indicate that NRMSEs are less than 0.06 regardless of the uncertainty conditions. Thus, it is reasonable to assume that the stochastic car-following distance follows a Gaussian distribution in a macroscopic statistical sense, and this assumption will be used for subsequent analysis.
@@
$\epsilon_{d}\sim$ $\mathcal{N}(0,0.1)$ $\mathcal{N}(0,0.5)$ $\mathcal{N}(0,1)$
$\epsilon_{\Delta v}\sim$ $\mathcal{N}(0,0.1)$ $\mathcal{N}(0,0.5)$ $\mathcal{N}(0,1)$ $\mathcal{N}(0,0.1)$ $\mathcal{N}(0,0.5)$ $\mathcal{N}(0,1)$ $\mathcal{N}(0,0.1)$ $\mathcal{N}(0,0.5)$ $\mathcal{N}(0,1)$
$\epsilon_{\dot{v}}\sim\mathcal{N}(0,0.1)$ 0.0170 0.0172 0.0208 0.0169 0.0168 0.0189 0.0166 0.0193 0.0198
$\epsilon_{\dot{v}}\sim\mathcal{N}(0,0.5)$ 0.0410 0.0278 0.0226 0.0408 0.0280 0.0236 0.0392 0.0280 0.0237
$\epsilon_{\dot{v}}\sim\mathcal{N}(0,1)$ 0.0590 0.0479 0.0408 0.0585 0.0478 0.0392 0.0579 0.0465 0.0388
@@ 3.3 Empirical AV data analysis
Considering that simulation experiments may not fully capture realistic driving behaviors, we also utilized real-world car-following data from commercial automated vehicles to validate their stochastic characteristics in longitudinal movements and verify our modified IDM in describing AV stochastic motion caused by robotic uncertainties. The Waymo Open Dataset (Ettinger et al., 2021) was selected as the primary dataset for the empirical analysis. Zhou et al. (2024) processed the dataset, focusing on longitudinal car-following scenarios with a constant-speed leading vehicle (LV) to accurately reflect the intricate behaviors of AVs, distinguishing it from the stochasticity caused by LV’s dynamics. Key data representing the states of the LV and the following AV (FAV), including spatial gaps, speeds, accelerations, and positions, are recorded. Our previous study (Zhou et al., 2024) identified that Waymo’s data collection methods make it the only dataset capable of effectively capturing speed fluctuations in AV operations while maintaining stability. The OpenACC dataset Makridis et al. (2021) with multiple car-following AVs, however, exhibits string instability, and the use of smoothing postprocessing methods to makes it unable to capture subtle stochastic motions. Therefore, considering that we isolated the impact of robotic uncertainty from LV dynamics on FAV motion through data filtering, it is well-suited for this study in identifying reliable results of stochastic car-following distances.
To identify stable car-following scenarios from complex real-world AV driving data, we conducted a sequence of data filtering steps. First, based on our previous work (Zhou et al., 2024) extracting unified car-following data, we selected scenarios where the preceding vehicle of the FAV maintained a constant speed, while the FAV tried to keep a consistent distance. Specifically, we retain the cases where the LV maintains a steady speed of around $20.2m/s$, the most common speed observed for the LVs, with a deviation of less than $0.2m/s$. Next, we excluded cases where the speed difference between two vehicles exceeded $1$ m/s, as larger differences may indicate an unstable car-following state, such as a free-flow approach or continuous braking. Finally, since the default car-following distance setting could vary across driving scenarios, we normalized the car-following distance in all sampled data to have a uniform variance of one and a mean of zero, ensuring comparability across the sub-datasets.
@@
Figure 4: A real car-following data result from Waymo dataset: (a) Acceleration of FAV (red dotted line) and distance between two vehicles (blue circled line) over time; (b) Histogram of distance between two vehicles (blue bars) and the fitted Gaussian distribution (red line).